OpenStudy (anonymous):

Factor: -3x^2+15xy-18y^2

OpenStudy (anonymous):

dam this ish

OpenStudy (anonymous):

-3(x^2-5xy+6y^2)=-3(x-2y)(x-3y)=3(2y-x)(x-3y)

OpenStudy (anonymous):

dam very nioce